5. A coin is tossed 400 times. Use the normal curve approximation to find the probability of obtaining


(a) between 185 and 210 heads inclusive;


(b) exactly 205 heads;


(c) fewer than 176 or more than 227 heads



1
Expert's answer
2022-06-06T08:34:04-0400

The probability of getting a head when a coin is tossed 1 time: "p=0.5."

Let's find the mean:

"\mu=np=400\cdot0.5=200."

The standard deviation:

"\sigma=\sqrt{np(1-p)}=\sqrt{400\cdot 0.5 \cdot (1-0.5)}=\sqrt{100}=10."


(a) The probability of obtaining between 185 and 210 heads inclusive:

"P(185\leq X \leq 210)=P(184.5 < X < 210.5)."

Let's find z-scores of 184.5 and 210.5:

"z_1=\frac{184.5-200}{10}=\frac{-15.5}{10}=-1.55,"

"z_2=\frac{210.5-200}{10}=\frac{10.5}{10}=1.05."

Now we have to use z-table.

"P(184.5 < X < 210.5)=P(-1.55<z<1.05)=0.8531-0.0606=0.7925."


(b) The probability of obtaining exactly 205 heads:

"P(X=205)=P(204.5<x<205.5)."

Let's find z-scores of 204.5 and 205.5:

"z_1=\frac{204.5-200}{10}=\frac{4.5}{10}=0.45,"

"z_2=\frac{205.5-200}{10}=\frac{5.5}{10}=0.55."

Now we have to use z-table.

"P(204.5<x<205.5)=P(0.45<z<0.55)=0.7088-0.6736=0.0352."


(c) The probability of obtaining fewer than 176 or more than 227 heads:

"P(X<176 \: or \: x>227)=P(X<175.5 \: or \: x>227.5)="

"=P(X<175.5)+P(x>227.5)."

Let's find z-scores of 175.5 and 227.5:

"z_1=\frac{175.5-200}{10}=\frac{-24.5}{10}=-2.45,"

"z_2=\frac{227.5-200}{10}=\frac{27.5}{10}=2.75."

Then we have:

"P(X<175.5)+P(x>227.5)=P(z<-2.45)+P(z>2.75)="

"=0.0071+(1-0.9970)=0.0071+0.003=0.0101."


Answer: (a) 0.7925 (b) 0.0352 (c) 0.0101